The American Baker’s Association reports that annual sales of bakery goods last year rose 15 percent, driven by a 50 percent increase in the demand for bran muffins. Most of the increase was attributed to a report that diets rich in bran help prevent certain types of cancer. You are the manager of a bakery that produces and packages gourmet bran muffins, and you currently sell bran muffins in packages of three. However, as a result of this new report, a typical consumer’s inverse demand for your bran muffins is now P = 11 - 1.5Q.

If your cost of producing bran muffins is C(Q) = 3.5Q, determine the optimal number of bran muffins to sell in a single package and the optimal package price.

Instruction: Enter your response for the optimal package price rounded to two decimal places.
Answer #1

MR 1-3Q - 3Q = 7.5 P = 11-1.5Q » TR = 110 - 1:54 C@)- 3.5Q - MC - 3.5 At profit maximizing level, MC = MR -> 3.5 =))-3Q - Q =